# VSF: Simple, Efficient, and Effective Negative Guidance in Few-Step Image Generation Models By Value Sign Flip

### v2

@@ Abstract
We introduce Value Sign Flip (VSF), a simple and efficient method for incorporating negative prompt guidance in few-step diffusion and flow-matching image generation models. Unlike existing approaches such as classifier-free guidance (CFG), NASA, and NAG, VSF dynamically suppresses undesired content by flipping the sign of attention values from negative prompts. Our method requires only small computational overhead and integrates effectively with MMDiT-style architectures such as Stable Diffusion 3.5 Turbo, as well as cross-attention-based models like Wan. We validate VSF on challenging datasets with complex prompt pairs and demonstrate superior performance in both static image and video generation tasks. Experimental results show that VSF significantly improves negative prompt adherence compared to prior methods in few-step models, and even CFG in non-few-step models, while maintaining competitive image quality. Code and ComfyUI node are available in https://github.com/weathon/VSF/tree/main.
Figure 1: Original image without negative guidance and image generated using our VSF negative guidance on Stable Diffusion 3.5 Large Turbo. The green prompt is the positive prompt, and the red one is the negative prompt. These examples have significant changes as they are removing essential parts of an object.
@@ 1 Introduction
Diffusion models (including flow matching models) have demonstrated their ability to produce diverse and high-quality images (Black Forest Lab, 2025; Woolf, 2022; sd, 3) and videos (Wan Team et al., 2025; Yin et al., 2025). However, a longstanding issue persists: negative guidance in image and video generation. Addressing this problem is crucial for improving content control, moderation (Schramowski et al., 2023), quality assurance, and reducing biases when generating general concepts (Chen et al., 2025a). However, vision language models (VLMs) have difficulties interpreting negations (Park et al., 2025; Alhamoud et al., 2025; Singh et al., 2025; 2024), rendering prompts containing negations ineffective (e.g., a prompt like “a scientist that is not wearing glasses” will usually generate a scientist with glasses, even more frequently than a plain “a scientist”). Classifier-free guidance (CFG) (Ho & Salimans, 2022) can address this issue when substituting unconditional generation with negative guidance.
@@
Figure 2: An example of when CFG is forcefully applied to step distilled models, the example is shown using a guidance scale of 2.8 and a step of 4 on SD-3.5-Large-turbo. The positive prompt is about a Canadian winter and a capybara, and the negative prompt is “snow”, we can see that it merges the two concepts unnaturally together and has severe over-saturation artifacts.
@@
To enhance efficiency in image and video generation, numerous models have been distilled to support inference in just a few steps (1-8 steps), such as Flux Schnell (Black Forest Lab, 2025), Stable Diffusion 3.5 Large Turbo (sd, 3), SDXL Lighting Lin et al. (2024), and CaucVid LoRA (Seppanen, ; Yin et al., 2025). However, CFG is incompatible with these models. These models are usually distilled and run in a guidance scale of 0 or 1 (depending on frameworks), which means only the positive guidance is used, and there is no extrapolation. When CFG is forcibly applied, the image will usually be over-saturated when the CFG scale is large enough to remove the unwanted concepts. Additionally, when the step counts are too small, the output shows both positive and negative prompts instead of explicitly avoiding the negative prompt (Nguyen et al., 2024) due to divergence between positive and negative guidance signals (Chen et al., 2025a). An example is shown in Figure 2. Additionally, even if CFG works, it requires two forward passes, one for positive guidance and one for negative guidance, which doubles the run time.
@@
To address this, two methods, NASA (Nguyen et al., 2024) and NAG (Chen et al., 2025a), have been introduced, employing negative guidance within attention space rather than the output space. NASA is currently limited to cross-attention models, while NAG primarily targets quality control rather than negative prompt avoidance. Both methods calculate positive and negative attentions separately and subtract them using a prefixed scale, resulting in a fixed guidance strength throughout the generation and on different areas on the image. This approach lacks adaptability to various time steps, layers, or image regions, limiting effectiveness in negative prompt adherence (Schramowski et al., 2023; Koulischer et al., 2025a; Ban et al., 2024).
@@
In this study, we introduce Value Sign Flip (VSF), a method that dynamically adjusts the guidance strength by flipping the sign of negative prompt values during attention. This enables the model to steer away from negative concepts adaptively based on their current presence strength, similar to the approach of Koulischer et al. (2025a). VSF has a small computational overhead and, when combined with few-step models, facilitates extremely fast image or video generation.
@@ 2 Related Work
2.1 Negation in Vision Language Models
Much previous work has shown that existing vision language models (VLM) struggle to understand negation (Yuksekgonul et al., 2023; Singh et al., 2024; Alhamoud et al., 2025; Park et al., 2025). In classification tasks, the model cannot correctly understand text with negation in it (e.g. “a dog running” vs “a dog not running” might have very close embeddings, even though they are opposite). This problem has been introduced into text-to-image generation tasks, making it hard for the model to generate images without certain concepts (examples in Figure 1 of Singh et al. (2025) and Figure 5 of Park et al. (2025)). Thus, classifier-free guidance (CFG) was used to introduce a negative prompt to the image generation process. More details in the next subsection. Several studies have attempted to tackle this issue by employing alternative training strategies, such as incorporating harder samples in the training data designed for negation tasks (Yuksekgonul et al., 2023; Singh et al., 2024; 2025; Alhamoud et al., 2025; Park et al., 2025). Some of these methods have shown improvements in image generation tasks. For instance, Park et al. (2025) reported gains in Neg Score—measuring whether the model retains the primary subject while correctly omitting the negated object—for both SD-1.4 and SDXL-1.0, by replacing the default CLIP encoder with their NegationCLIP on their dataset, without additional T2I training. Nonetheless, the Neg Score remained below 0.5, indicating limited effectiveness.
These methods generally require re-training the text encoder (usually a CLIP-like model) with contrastive learning, which poses challenges for models that do not use contrastively pre-trained encoders, such as T5 (Raffel et al., 2023) in Stable Diffusion 3 (sd, 3; Esser et al., 2024) and Flux (Black Forest Lab, 2025). Moreover, each model using a different text encoder would require a separate, dedicated adaptation.
@@ 2.2 Classifier Free Guidance
Original classifier-free guidance (CFG) (Ho & Salimans, 2022) generates a conditioned noise prediction and an unconditioned noise prediction. In flow matching (Lipman et al., 2023), the predicted targets are the velocity ($u_{t}$) pointing to the image. Thus, the original flow matching CFG prediction can be written as
$u_{t}=f(\emptyset,x_{t+1})+\lambda(f(p^{+},x_{t+1})-f(\emptyset,x_{t-1})),$ (1)
@@
$u_{t}=f(p^{-},x_{t+1})+\lambda(f(p^{+},x_{t+1})-f(p^{-},x_{t+1})),$ (2)
2.3 Recent Works on Negative Guidance
The studies on negative guidance are very limited (only (Ban et al., 2024; Koulischer et al., 2025a; Schramowski et al., 2023)). Ban et al. (Ban et al., 2024) finds that the negative prompts affect the model by delayed effects and neutralization. After the model has generated unwanted contents, the negative guided vector ($u_{p^{-}}$) will neutralize the content. They also observed the reverse activation effect, where the negative prompt introduced early in the diffusion processes could actually induce the unwanted concepts. To address this, they proposed applying the negative guidance later in the diffusion process and found it effective.
Schramowski et al. (2023) used a very similar idea as CFG to avoid unwanted (NSFW) content. They generate an unsafe vector and purposely avoid it by subtracting it from the predicted noise. They also added a pixel-level guidance scale that depends on the pixel-wise distance between the positive predicted noise and the unwanted noise.
Koulischer et al. (2025a) used similar ideas of both and proposed a temporal dynamic guidance scale method. They calculate a probability that the generated concept contains negative content and adjust the guidance scale accordingly. However, their adaptive scale only changes throughout the steps and does not adapt to different regions in the image.
After completing the manuscript, we found a public GitHub repository for a cross-attention module that might use a similar method to ours; however, it provides no explanation, documentation, paper, or tests, so its performance and exact approach cannot be verified  [^1] https://github.com/hako-mikan/sd-webui-negpip.
@@ 2.4 Few-Step Image Generation Models
Traditional diffusion or flow-matching image generation models typically require many inference steps. However, with improved schedulers, this can be reduced to around 20 steps. Recent approaches go further by using step distillation to reduce the number of steps to fewer than 8, or even a single step, as demonstrated in Flex Schnell (Black Forest Lab, 2025), SDXL Lightning (Lin et al., 2024), CausVid (Seppanen, ; Yin et al., 2025), and Stable Diffusion 3.5 Turbo (sd, 3). Since these models are distilled, they generally do not use classifier-free guidance (CFG) during inference; when CFG is forcibly applied, the results are significantly degraded to the point that it is completely unusable (Nguyen et al., 2024), see Figure 2 for an example.
@@ 2.5 Recent Works on Negative Guidance in Few-Step Models
Recently, two approaches have specifically targeted negative guidance techniques for few-shot models: Negative-Away Steer Attention (NASA) (Nguyen et al., 2024) and Normalized Attention Guidance (NAG) (Chen et al., 2025a). Although they both focused on avoiding unwanted content and improving quality (using a negative prompt describing bad quality), NASA mainly focused on avoiding unwanted content, while NAG focused on improving quality.
The authors of the NASA study found that neither standard CFG nor CFG applied directly to text embeddings yields desirable results in few-step scenarios, particularly in single-step settings. Specifically, the regular CFG independently computes positive and negative guidance signals, preventing the negative guidance from effectively neutralizing unwanted concepts. As a result, the produced images merely appear as a mixture of both positive and negative prompts unnaturally (an average image of the positive prompt generated image and the negative prompt independently generated image) rather than excluding negative prompt elements. Furthermore, the authors noted that applying CFG to text embeddings produces minimal benefits. For detailed examples and further illustration, readers are referred to the original paper introducing NASA (SNOOPI (Nguyen et al., 2024)).
The method NASA proposed is to apply the guidance in intermediate states instead of the predicted noise or velocity. Specifically, they calculate a positive attention output $Z^{+}$ and a negative attention output $Z^{-}$, and they output the final attention $Z^{NASA}$ by subtracting the two with a factor $\alpha$, as shown in Equation 3. The alpha value is usually between 0 and 1.
@@
3 Proposed Methods
Figure 3: The attention mechanism of our method. We pass in image tokens (II), positive prompt tokens ($P$), and negative prompt tokens (NN) into attention. For key and values, NN is duplicated, with values of one copy ($N^{(1)}$ scaled by $\alpha$. Some areas are masked to avoid interference. An bias $\beta$ is added to $I\rightarrow N^{(1)}$ attention.
@@ 3 Proposed Methods
However, to maintain the stability of the attention space, they also applied normalization to $\widetilde{Z}^{NAG}$ to limit its norm releative to $Z^{+}$ with scale $\tau$, resulting in $\hat{Z}$. Then it used a blending factor $\alpha$ to blend it with the positive attention result, as shown in Equation 5.
@@
The normalization and blending ensure the attention output of the NAG does not drift away from what the model usually sees during training, improving the quality of generated images. However, it also limits the model’s following to negative prompt guidance if the constraint is set to be too tight (i.e., high $\alpha$ and low $\tau$).
@@ 3.1 Value Sign Flip Adaptive Attention
Our proposed method is similar to NASA. NASA used a fixed value $\alpha$ at all layers and stages for all tokens during the image generation, pushing the content away from the negative output ($Z^{-}$) even when the unwanted concept is not present. Previous work (Koulischer et al., 2025b; a; Ban et al., 2024; Schramowski et al., 2023) has shown that a dynamic or adaptive guidance scale could yield better results. In the NAG (Chen et al., 2025a) future work section, they also hypothesized that token-level modulation may be beneficial. Additionally, in NAG, they applied guidance by extrapolation, which means that to increase the negative guidance scale, you will also need to increase the positive guidance scale. This might make it challenging for cases where positive concepts are closely related to negative concepts (e.g., bike but no wheels) or when the effects needed are contradictory to extrapolation (e.g., when the need to generate camouflaged, undersaturated, or blurry samples is purposely done).
Drawing from Koulischer et al. (2025a) and Schramowski et al. (2023), one possible solution is to steer the latent space away when the model is about to generate, or has already generated, unwanted content. In their approach, a probability-based method is used to determine the appropriate guidance factor. Alternatively, a more intuitive method involves using the model’s attention map: when the image attends more to the negative prompt compared to the positive one, it should be steered away stronger accordingly. This can be implemented by simply concatenating the values and keys of the positive and negative prompts, then flipping the sign of the negative prompt values so that when the image attends to the negative prompt, the flipped value of the negative prompt can cancel the unwanted content. Note that the key of the negative prompt is not flipped to keep the original meaning of the unwanted concept to match image patches. The equation of our method in cross attention models, written in the matrix calculation, is shown in Equation 6, where $\oplus$ means matrix concatenation on the sequence length dimension, and $\sigma$ is the softmax function on the sequence length dimension, $Q$ is the image query tokens, $K^{+}$ and $K^{-}$ are the positive and negative prompt keys, $V^{+}$ and $V^{-}$ are the positive and negative prompt values, and $\alpha$ is the factor controlling the strength of the guidance.
@@
Mathematically, this is equivalent to computing the ratio between the image-to-positive and image-to-negative attention map strength, scaling each attention output using this ratio, and subtracting the negative velocity from the positive one.

### v5

@@ Abstract
We introduce Value Sign Flip (VSF), a simple and efficient method for incorporating negative prompt guidance in few-step (1-8 steps) diffusion and flow-matching image and video generation models. Unlike existing approaches such as classifier-free guidance (CFG), NASA, and NAG, VSF dynamically suppresses undesired content by flipping the sign of attention values from negative prompts. Our method requires only a small computational overhead and integrates effectively with MMDiT-style architectures such as Stable Diffusion 3.5 Turbo and Flux Schnell, as well as cross-attention-based models like Wan. We validate VSF on a challenging dataset, NegGenBench, with complex prompt pairs. Experimental results on our proposed dataset show that VSF significantly improves negative prompt adherence (reaching 0.420 negative score for quality settings and 0.545 for strong settings) compared to prior methods in few-step models (scored 0.320-0.380 negative score) and even CFG in non-few-step models (scored 0.300 negative score), while maintaining competitive image quality and positive prompt adherence. Our method is also a suppressed generate-then-edit pipeline, while also having a much faster runtime. Code, ComfyUI node, and dataset will be released.
Figure 1: Original image without negative guidance and image generated using our VSF negative guidance on Stable Diffusion 3.5 Large Turbo. The green prompt is the positive prompt, and the red one is the negative prompt. These examples have significant challenges as they are removing essential parts of an object. The “hands” in the last image mean clock hands.
@@ 1 Introduction
Diffusion models (including flow matching models) have demonstrated their ability to produce diverse and high-quality images (Black Forest Lab, 2025; Woolf, 2022; Stability AI, 2024) and videos (Wan Team et al., 2025; Yin et al., 2025). However, a longstanding issue remains: the challenge of effectively applying negative guidance in image and video generation. Addressing this problem is crucial for improving content control, moderation (Schramowski et al., 2023), quality assurance, and reducing biases when generating general concepts (Chen et al., 2025a). However, vision language models (VLMs) have difficulties interpreting negations (Park et al., 2025; Alhamoud et al., 2025; Singh et al., 2025; 2024), rendering prompts containing negations ineffectively or made the negative prompt appears even more (e.g., a prompt like “a scientist who is not wearing glasses” will often generate a scientist with glasses—sometimes even more frequently than a simple prompt like “a scientist”). Classifier-free guidance (CFG) (Ho & Salimans, 2022) can be used to address this issue when substituting unconditional generation with negative guidance.
@@
Figure 2: An example of forcefully applying CFG to a step-distilled model is shown using a guidance scale of 2.8 and only 4 steps on SD-3.5-Large Turbo. The positive prompt describes a Canadian winter with a capybara, while the negative prompt includes the word “snow.” The resulting image merges these conflicting concepts unnaturally and exhibits severe over-saturation artifacts.
@@
However, to enhance efficiency in image and video generation, numerous models have been distilled to support inference in just a few steps (1-8 steps), such as Flux Schnell (Black Forest Lab, 2025), Stable Diffusion 3.5 Large Turbo (Stability AI, 2024), SDXL Lighting Lin et al. (2024), SNOOPI (Nguyen et al., 2024), and CaucVid (Seppanen, ; Yin et al., 2025). However, CFG is incompatible with these models. These models are usually distilled and run in CFG-disabled mode, which means only the positive guidance is used, and there is no extrapolation. When CFG is applied forcefully, the resulting image often becomes oversaturated, particularly when the CFG scale is set high enough to suppress unwanted concepts. Moreover, if the number of diffusion steps is too low, the output may reflect features from both the positive and negative prompts (Nguyen et al., 2024), rather than excluding the negative prompt entirely. This occurs due to a divergence between the positive and negative guidance signals (Chen et al., 2025a). An example is shown in Figure 2. Additionally, even if CFG works, it requires two forward passes, one for positive guidance and one for negative guidance, which doubles the run time.
@@
To address this, two methods, Negative Steer Away Attention (NASA) (Nguyen et al., 2024) and Normalized Attention Guidance (NAG) (Chen et al., 2025a), have been introduced, employing negative guidance within attention final output space rather than the output space. NASA is currently limited to cross-attention models (though it can be re-implemented into other models), while NAG primarily targets quality control rather than negative prompt avoidance. Both methods calculate positive and negative attentions separately and subtract them using a prefixed scale (same as CFG), resulting in a fixed guidance strength throughout the generation, across different areas of the image, and at different layer of the model. This approach lacks adaptability to various time steps, layers, or image regions, limiting effectiveness in negative prompt adherence compared to a more adaptive method (Schramowski et al., 2023; Koulischer et al., 2025; Ban et al., 2024).
@@
In this study, we introduce Value Sign Flip (VSF), a method that dynamically adjusts the guidance strength by flipping the sign of negative prompt values within the attention calculation (i.e., not the attention output). This enables the model to steer away from negative concepts adaptively based on their current presence strength, similar to the approach of Koulischer et al. (2025). VSF has a small computational overhead and, when combined with few-step models, facilitates extremely fast image or video generation (¡3 seconds). Our contributions in this work are: (1) we proposed a new method for better negative guidance; (2) we constructed a dataset, NegGenBench, consisting of challenging positive-negative prompt pairs; (3) we collected images generated using these three methods (VSF, NAG, NASA) and labeled their negative and quality score. We further fine-tuned a VLM on it for future work to better evaluation of negative prompt following.
@@ 2.1 Classifier Free Guidance
Vision language models struggle to understand negation (Yuksekgonul et al., 2023; Singh et al., 2024; Alhamoud et al., 2025; Park et al., 2025) (We discussed more about this in the Appendix). Original classifier-free guidance (CFG) (Ho & Salimans, 2022) generates a conditioned noise prediction and an unconditioned noise prediction. In flow matching (Lipman et al., 2023), the predicted targets are the velocity ($u_{t}$). Thus, the original flow matching CFG prediction can be written as
$u_{t}=f(\emptyset,x_{t+1},t)+\lambda(f(p^{+},x_{t+1},t)-f(\emptyset,x_{t-1},t)),$ (1)
@@
$u_{t}=f(p^{-},x_{t+1},t)+\lambda(f(p^{+},x_{t+1},t)-f(p^{-},x_{t+1},t)),$ (2)
where $p^{-}$ is the negative prompt.
2.2 Recent Works on Dynamic Negative Guidance
The studies on dynamic negative guidance are very limited (only (Ban et al., 2024; Koulischer et al., 2025; Schramowski et al., 2023)). Ban et al. (Ban et al., 2024) found that the negative prompts affect the model by delayed effects and neutralization. After the model has generated unwanted contents, the negative guided output ($u_{p^{-}}$) will neutralize the content. They also observed the reverse activation effect, where the negative prompt introduced early in the diffusion processes could actually induce the unwanted concepts. To address this, they proposed applying the negative guidance later in the diffusion process and found it effective.
Schramowski et al. (2023) used a very similar idea as CFG to avoid unwanted (NSFW) content. They generate an unsafe vector and purposely avoid it by subtracting it from the predicted noise. They also added a pixel-level guidance scale that depends on the pixel-wise distance between the positive predicted noise and the unwanted noise, making it adaptive to different regions in the image.
Koulischer et al. (2025) used similar ideas of both and proposed a temporal dynamic guidance scale method. They calculate a probability that the generated concept contains negative content and adjust the guidance scale accordingly. However, their adaptive scale only changes throughout the steps and does not adapt to different regions in the image.
@@ 2.3 Few-Step Image Generation Models
Traditional diffusion or flow-matching image generation models typically require many inference steps. However, with improved sampler, this can be reduced to around 20 steps. Recent approaches go further by using step distillation to reduce the number of steps to fewer than 8, or even a single step, as demonstrated in Flex Schnell (Black Forest Lab, 2025), SDXL Lightning (Lin et al., 2024), CausVid (Seppanen, ; Yin et al., 2025), SNOOPI (Nguyen et al., 2024), and Stable Diffusion 3.5 Turbo (Stability AI, 2024). Since these models are distilled, they generally do not use classifier-free guidance (CFG) during inference; when CFG is forcibly applied, the results are significantly degraded to the point that it is completely unusable (Nguyen et al., 2024), see Figure 2 for an example.
@@ 2.4 Recent Works on Negative Guidance in Few-Step Models
Recently, two approaches have specifically targeted negative guidance techniques for few-shot models: Negative-Away Steer Attention (NASA) (Nguyen et al., 2024) and Normalized Attention Guidance (NAG) (Chen et al., 2025a). Although they both focused on avoiding unwanted content and improving quality (using a negative prompt that describes bad quality), NASA mainly focused on avoiding unwanted content, while NAG focused on improving quality.
The authors of the NASA study found that neither standard CFG nor CFG applied directly to text embeddings yields desirable results in few-step scenarios, particularly in single-step settings. Specifically, the regular CFG independently computes positive and negative guidance signals, preventing the negative guidance from effectively neutralizing unwanted concepts. As a result, the produced images merely appear as a mixture of both positive and negative prompts unnaturally (an average image of the positive prompt generated image and the negative prompt independently generated image) rather than excluding negative prompt elements. Furthermore, the authors noted that applying CFG to text embeddings produces minimal benefits. For detailed examples and further illustration, readers could refer to the original paper introducing NASA (Nguyen et al., 2024).
The method NASA applies the guidance in intermediate states instead of the predicted noise or velocity. Specifically, they calculate a positive attention output $Z^{+}$ and a negative attention output $Z^{-}$, and the final attention $Z^{NASA}$ is obtained by subtracting the two with a factor $\alpha$, as shown in Equation 3. The alpha value is usually between 0 and 1.
@@
Figure 3: The attention mechanism of our method. We pass in image tokens (II), positive prompt tokens ($P$), and negative prompt tokens (NN) into attention. For key and values, NN is duplicated, with values of one copy ($N^{(1)}$) scaled by $\alpha$. Some areas are masked to avoid interference. An bias $\beta$ is added to $I\rightarrow N^{(1)}$ attention.
@@
However, to maintain the stability of the attention output space, they also applied normalization to $\widetilde{Z}^{NAG}$ to limit its norm relative to $Z^{+}$ with scale $\tau$ per token, resulting in $\hat{Z}$. Then it used a blending factor $\alpha$ to blend it with the positive attention result, as shown in Equation 5.
@@
The normalization and blending ensure the attention output of the NAG does not drift away from what the model usually sees during training, improving the quality of generated images. However, it also limits the model’s ability to follow to negative prompt guidance if the constraint is set to be too tight (i.e., high $\alpha$ and low $\tau$).
3 Proposed Methods
Our proposed method is built on top of NASA (Nguyen et al., 2024), Koulischer et al. (2025), and Schramowski et al. (2023). NASA has a fixed guidance for every attention calculation, Koulischer et al. (2025) does not have token-level modulation, and Schramowski et al. (2023)’s approach of taking the differences between the positive and negative noise predictions is too simple for complex situations, as mentioned in Koulischer et al. (2025). In NAG’s (Chen et al., 2025a) future work section, they also mentioned the possibility of token-level modulation but they did not propose a specific solution.
@@ 3.1 Value Sign Flip Adaptive Attention
We propose to expand Koulischer et al. (2025) idea to token-level modulation in few-step models. Let $W$ be a per-token weight at each attention calculation for how strongly the token is associated with a positive concept compared to the negative one. We can modify the NASA attention to Eq. 7. $W$ is obtained by a function with a positive prompt, a negative prompt, and an image as input.
$W=g(p^{+},p^{-},I),$ (6)
then we can rewrite the equation in NASA as
$Z^{W}=WZ^{+}-\alpha(1-W)Z^{-}$ (7)
An intuitive method to calculate $W$ involves using the model’s attention map: when the image attends more to the negative prompt compared to the positive one, it should be steered away strongly accordingly. Thus, we can calculate the attention map between the image and the positive tokens $A^{+}$ and the image and the negative tokens $A^{-}$ before softmax calculation, then calculate their ratios to their sum. $Q$ is the image query tokens and $K^{+}$ and $K^{-}$ are the positive and negative prompt keys.
$A^{+}=\exp(\frac{Q(K^{+})^{T}}{\sqrt{d}}),A^{-}=\exp(\frac{Q(K^{-})^{T}}{\sqrt{d}}),W=\frac{\sum A^{+}}{\sum(A^{+}+A^{-})}$ (8)
This approach involves complex attention calculation and two attention passes, but it can be implemented by a simpler approach. We can concatenate the values and keys of the positive and negative prompts, then flip the sign of the negative prompt values. This enables that when the image attends to the negative prompt, the flipped value of the negative prompt can cancel the unwanted content. The equation of our method in cross attention models, written in the matrix calculation, is shown in Equation 9, where $\oplus$ means matrix concatenation on the sequence length dimension, $\sigma$ is the softmax function on the sequence length dimension, and $V^{+}$ and $V^{-}$ are the positive and negative prompt values.
@@
This is similar to noise-canceling headphones, where a “flipped” wave is played to cancel the noise. Note that the key of the negative prompt is not flipped to keep the original meaning of the unwanted concept to match image patches. Mathematically, this is equivalent to $Z^{W}$. Proof is in the Appendix.